# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

### v1

@@ Abstract
Vision Language Models (VLMs) can be trained more efficiently if training sets can be reduced in size. Recent work has shown the benefits of masking text during VLM training using a variety of approaches: truncation, random masking, block masking and syntax masking. In this paper, we show that the best masking strategy changes over training epochs and that, given sufficient training epochs, word frequency information is what you need to achieve the best performance. Experiments on a large range of data sets demonstrate the advantages of our approach, called Contrastive Language-Image Pre-training with word Frequency Masking (CLIPF). The benefits are particularly evident as the number of input tokens decreases. We analyze the impact of CLIPF vs. other masking approaches on word frequency balance and discuss the apparently critical contribution of CLIPF in maintaining word frequency balance across POS categories.
@@ 1 Introduction
This paper contributes to the recent trend of text-masking approaches that improve VLMs by removing words from the text data during training, speeding up training in the process [11, 12]. We introduce CLIPF (Contrastive Language-Image Pre-training with word Frequency Masking), which selects words to be masked on the basis of word frequency. We carry out a comparison with CLIPA masking [12], which selects words on the basis of four strategies that do not use frequency information: truncation, random word selection, random block selection, and part-of-speech (POS) information and is currently considered the state-of-the-art text-masking.
The comparison of the learning curves of our CLIPF with CLIPA masking in Fig. 1 illustrates two key contributions of this paper. First, we demonstrate that the number of training epochs is crucial to which masking strategy yields the highest performance. The finding is an important improvement on the state of the art of text masking for CLIP training, represented by [12]. In [12], CLIP is trained on a limited number of epochs and syntax masking is found to outperform truncation, random, and block masking. In contrast, we show that random and block masking outperform syntax masking after a sufficient number of epochs. Second, we show that when masking text for CLIP training, frequency is what you need. As evident in Fig. 1, our frequency-based CLIPF approach outperforms all CLIPA masking approaches with sufficient epochs of training, leading it to be the best performing text masking approach.
Table 1: Illustration of different text masking methods and words masking probabilities. Bold percentages or check marks provide an example of the words that would be retained by each text masking strategy. For the Truncation and Syntax methods, four words per text are retained as the input. For the Random, Block, and CLIPF methods, we present the word masking probabilities. The second row shows the part-of-speech (POS) category of each word.
Models Methods Words
a small dog sleeping on a couch next to a remote .
POS Other Adjective Noun Verb Other Other Noun Other Other Other Noun Other
CLIPA Truncation
Random 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50%
Block 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50%
Syntax
CLIPF Frequency 99.37% 96.28% 96.13% 88.35% 98.96% 99.37% 87.03% 93.69% 98.89% 99.37% 81.11% 99.37%
@@
Figure 1: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies. The text masking ratio is 75%. The backbone of the image encoder is ViT-B/16. The model is pre-trained on CC12M for 30 epochs. We use image masking (75%) to speed up training. All approaches are subsequently fine-tuned with one epoch on the full, unmasked data.
@@
The history of text masking in language model training dates to [13], where subsampling of frequent words in the training text was used to improve the performance of skip-gram language models. The study of text masking for VLMs got off to a slow start, since initially [8] reported that text masking, although it promotes efficiency, deteriorated performance. Subsequently, [12] expanded the types of masking studied and demonstrated that masking can indeed improve performance. Simultaneously, frequency-masking was proposed at a workshop [11], but fell short of the potential of frequency realized by CLIPF due to the use of a cut-off threshold. To our knowledge, our work reported in this paper is the first to attempt to discover the reasons that certain forms of text masking outperform others.
An initial illustration of the differences between task-masking strategies is provided in Table 1. Truncation consistently retains the beginning of the text. With the random (random word selection) and block (random block selection) methods, each word has an equal probability of being masked, with block masking maintaining word neighborhoods. Syntax masking prioritizes the retention of nouns, followed by other part-of-speech (POS) categories. Masking based on frequency (CLIPF) has the potential to address the shortcomings of the other approaches. Unlike truncation, it does not prioritize words that tend to occur at the beginning of the sentence. In contrast to random and block masking, it represses words with the POS category “Other”, which are often stop words and for this reason high frequency. However, it does not give an overly high priority to nouns and adjectives, as with syntax masking.
In sum, the contributions of this paper are as follows:
@@
  • We point out, to our knowledge for the first time, the importance of the number of training epochs for comparing text masking approaches.
  • We introduce CLIPF, a frequency-based masking approach that outperforms previously proposed approaches.
  • We carry out an analysis that provides insight into why CLIPF outperforms other approaches.
@@
Our work is reproducible because it uses training data sets that are openly available and small enough to allow researchers with limited computational resources to repeat our experiments and because we release our code [^1] https://github.com/MingliangLiang3/CLIPF
@@ 2.1 Vision-Language Models
Vision-language models (VLMs), such as CLIP, ALIGN, LiT, and FLIP [3, 4, 10, 8], learn image embeddings from large-scale natural language supervision using contrastive learning [14] to grants the models remarkable transferability. Because CLIP does not release its pre-training dataset, OpenCLIP pre-trained models on several large-scale public datasets [15, 16, 17, 18] to reproduce CLIP. This implementation adheres to the scaling law observed in Neural Language Models [19] and Vision-Language Models [3], where loss decreases according to a power-law relationship with model size, dataset size, and computational resources utilized during training. While large language models (LLMs) have advanced rapidly, we focus on vision-language models (VLMs) like CLIP, as they are essential tools for bridging visual and textual modalities in current research. Although CLIP and OpenCLIP are pre-trained on very large-scale datasets, recent works like DeCLIP [20] and SLIP [5] demonstrate that Vision-language Models (VLMs) can be trained efficiently with smaller datasets, enabling researchers with limited computational resources to contribute. Moreover, masking image patches or reducing image resolution during VLM pre-training also substantially decreases computational costs while maintaining or enhancing performance, as shown by techniques like FLIP, RECLIP, CLIPA [8, 21, 12].
@@ 2.2 Text masking for VLM training
Masking text presents another viable option to speed up the pre-training of VLMs, and is particularly effective alongside strategies that employ a high ratio of image patch masking. FLIP further explored text masking techniques, namely, random and truncation masking, to enhance efficiency [8]. While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retrains the words that preceded them.
Subsequently, CLIPA [12] introduces block masking and syntax masking. As already mentioned in Section 1, CLIPA [12] shows text masking during training can improve the performance of a VLM. CLIPA syntax masking prioritizes the retention of nouns, followed by adjectives and then other words. In [12], syntax masking outperforms the truncation, random, and block approaches. However, [12] does not vary the number of training epochs. An important contrast with our work, as previously mentioned, is our attention to the number of training epochs. Further, CLIPF, as we will see, counters the danger of giving overly high priority to nouns, leading to overfitting, and also avoids overly suppressing words of the POS category “Other”.
An initial effort to exploit frequency information for text masking during VLM training was made by SW-CLIP [11]. Unfortunately, SW-CLIPF introduces a frequency cutoff threshold, which reduces performance, as we demonstrate later in this paper. In order to exploit frequency information effectively each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
3 Method
In this section, we describe various text masking strategies, including our CLIPF method. All these strategies accelerate pre-training by masking specific words in the text.
3.1 Text Masking
@@
This function $f$ determines which words are masked to get the modified text $T_{new\_i}$. And, we define the number of input tokens as $k$. We will sequentially introduce our masking strategies: random, truncation, block, syntax, and frequency. It is important to note that the number of padding tokens remains consistent across all these strategies.
@@
3.2 Word Frequency Masking
As in previous work, subsampling of frequent words was introduced by [13] in the natural language processing (NLP) domain to accelerate the learning process for word representations. The goals are to accelerate training and to address the imbalance between frequent and infrequent words in the training corpus. In our work, we use the formula (Formula 2) introduced in [13] to determine the probability of masking a word during training. Thus, the complete formula to calculate the probability $P(w_{i})$ that accounts for all scenarios is:
$P(w_{i})=\begin{cases}1&\text{if}N<5\\ 0&\text{if}f(w_{i})<t\\ 1-\sqrt{\frac{t}{f(w_{i})}}&\text{otherwise}\end{cases}$ (2)
Here, the masking probability of the word $w_{i}$ is $P(w_{i})$, $t$ is the threshold, and $f(w_{i})$ is the frequency of the word. Additionally, if the total number of words ($N$) is fewer than 5, deemed too infrequent for effective learning, the probability that the word will be masked is 1.
Following FLIP [8], we fine-tune the model with an additional epoch to eliminate the distribution gap between the masking and unmasking.
@@ 4 Experimental Setup
Our implementation follows CLIP [3], OpenCLIP [9], FLIP [8] and CLIPA [12]. In this section, we present the details of the training and testing.
4.1 Implementation
Dataset We pre-train the models on Conceptual Captions 3M (CC3M) [23] and Conceptual 12M (CC12M) [15]. These datasets were selected due to their diverse range of real-world concepts and scenes. Popular VLMs such as OpenCLIP, DeCLIP, and SLIP [9, 20, 5] also employ these datasets for training, leveraging their extensive and diverse content. Due to the expiration and inaccessibility of some URLs in the training set, we have downloaded 2.7 million image-text pairs for CC3M and 9.3 million image-text pairs from CC12M. The characteristics of the dataset are presented in Table 2. The CC12M dataset is 3.42 times larger than the CC3M dataset, with captions that are on average 2.15 times longer. Since the number of input tokens is fixed, the impact of different text masking ratios varies across datasets due to differences related to caption length.
@@ 4.1 Implementation
Architecture For the image encoder, we use ViT-B/16 [24] architectures with global average pooling and learnable positional embeddings. The input image size is 224. For the text encoder, we use a Transformer-based model [25] and byte-pair encoding with a 49K token vocabulary. We chose these encoders because they are the largest encoders that we can use given our resources. When using full text for training, the maximum context length is 32. Zero-padding is applied to input text that is shorter than the maximum token length of the model.
We trained the models using 8 RTX A5000 GPUs with the same settings to ensure consistent conditions across all models. This allows us to effectively measure and compare the performance of each model in processing text tokens during training. We experiment with different text token input sizes, namely, 32, 16, 8, 6, and 4 text tokens. As the input size gets smaller, we can increase the batch size, maximizing computational memory usage. We used batch sizes of 664, 832, 896, 928, and 944, respectively. Note that CLIP (and therefore CLIPA and CLIPF) is trained using contrastive learning, which benefits from larger batch sizes. Note also that although we used different text masking with the same settings and batch sizes, syntax masking was slower than the other text masking strategies when conducting the experiments because POS processing for each word is time-consuming.
@@
augmentation RandomResizedCrop RandomResizedCrop
Table 4: During pre-training, each model processes a certain number of image and text tokens. The first and second columns of the table display the image and text masking ratios, respectively. The third and fourth columns show the number of image and text tokens processed after masking. The fifth column presents the total number of tokens (image + text) after masking. The last column provides the percentage of total tokens relative to the unmasked case.
@@
Training We follow the setup of OpenCLIP [9], FLIP [8] and CLIPA to pre-train our model. Following OpenCLIP [9], we pre-train our model for 30 epochs on the CC3M and CC12M datasets. Details of the pre-training configuration are given in Table  3.
@@
During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4. To measure the training speed of CLIPF, CLIP, and CLIPA, we compare the number of text tokens processed by each model. As shown in Table 4, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios. When we pre-trained the model using image masking, the total number of tokens is 81 and the percentage of text tokens is 39.5%. When we apply 50% text masking, the total number of tokens is 65. Compared to training without text masking, this results in a speed increase of approximately 20%. Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
Fine-tuning Following FLIP, and CLIPA, in order to bridge the gap between pre-training and evaluation, we fine-tuned the model without images and text masking. The details of the fine-tuning configuration are provided in Table 3.
Table 5: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K. reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine tuned for an additional epoch on all data (fine-tune). “Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
@@ 4.2 Evaluation
Evaluation setup Following the CLIP [3], FLIP [8], and CLIPA [12] methods, several classification benchmarks were used to evaluate the zero-shot capability of CLIPF. Among these benchmarks is ImageNet-1K, a widely recognized dataset in computer vision. It is frequently used for image classification and VLMs tasks and comprises 50K validation samples across 1K different classes. We filled each class into the templates provided by CLIP [3] to calculate the average of the text embeddings. We use the same evaluation settings as CLIP [3] to evaluate the other downstream datasets.

### v2

@@ Abstract
Vision Language Models (VLMs) can be trained more efficiently if training sets can be reduced in size. Recent work has shown the benefits of masking text during VLM training using a variety of approaches: truncation, random masking, block masking and syntax masking. In this paper, we show that the best masking strategy changes over training epochs and that, given sufficient training epochs. We analyze existing text masking approaches including syntax masking, which is currently the state of the art, and identify the word frequency distribution as important in determining their success. Experiments on a large range of data sets demonstrate that syntax masking is outperformed by other approaches, given sufficient epochs, and that our proposed frequency-based approach, called Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF) has numerous advantages. The benefits are particularly evident as the number of input tokens decreases.
@@ 1 Introduction
However, pre-training CLIP models typically requires thousands of GPU-days [3, 9]. To accelerate the pre-training process, FLIP, A-CLIP, and CLIPA use masking strategies for image and text tokens when pre-training CLIP models [8, 11, 12]. In addition to the speed-up, masking makes it possible use large batch sizes, which enhances the benefits of contrastive learning and has been considered as a form of regularization, reducing the risk of overfitting [8]. For these reasons, despite the loss of some information from the original image and text, masking strategies have the potential to improve model performance.
@@
Figure 1: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16. The models are pre-trained on CC12M for 45 epochs with 75% text masking ratio. We also use image masking (75%) to speed up pre-training. All approaches are subsequently fine-tuned with one epoch on the full, unmasked data.
Currently, the state of the art in text masking is represented by CLIPA [12], which compares four masking strategies: truncation, random word selection, random block selection, and syntax-based selection, i.e., masking on the basis of part-of-speech (POS) information. The conclusion is that syntax masking is the strategy that, overall, provides the best quality. In this paper, we carry out a set of experiments that demonstrate that this conclusion must be revisited. Specifically, as shown in Fig. 1, if enough epochs are used, other forms of masking substantially outperform syntax masking. The conclusion of CLIPA [12] misses the weakness of syntax masking because it trains with a limited number of epochs (6.4 epochs).
The special focus of our paper is a frequency masking approach that we introduce: CLIPF (CLIP Frequency masking). In Fig. 1, we see that with enough epochs CLIPF and CLIPA block masking perform similarly. We provide an analysis that demonstrates why CLIPF, block masking, and random masking emerge as clearly superior to syntax masking and truncation. The analysis studies the distribution of words by frequency of occurrence in the training data after masking. It shows that good masking strategies maintain helpful properties of this distribution and that syntax masking distorts the distribution of words over POS categories. These observation leads to the main insight of our paper, namely, when masking text for VLM pre-training, what you need for best results is to pay attention to the impact of masking on word frequency.
Our CLIPF approach is useful in practice because it performs relatively well in the range of 16-40 epochs. This range is a typical number of epochs to train a system in practice. For example, Open_CLIP [9] trains on 30 epochs. In Fig. 1, we see that in this range CLIPF outperforms other masking strategies. We carry out a set of experiments with 30 epochs to shed further light on the role of frequency in masking by comparing CLIPF vs. the different variants of CLIPA masking on zero-shot tasks and on image retrieval. We also investigate training on a larger data set.
@@
The history of text masking in language model training dates to [13], where subsampling of frequent words in the training text was used to improve the performance of skip-gram language models. The study of text masking for VLMs got off to a slow start, since initially [8] reported that text masking, although it promotes efficiency, deteriorated performance. The CLIPA [12] work is important because it expanded the types of masking studied and demonstrated that masking can indeed improve VLM performance. To our knowledge, no previous work has attempted to discover the reason that certain forms of text masking outperform others or identified the importance of the way in which text masking approaches impact word frequency.
In sum, our paper makes the following contributions to the study of text masking in pre-training VLMs:
@@
  • We point out the importance of the number of training epochs when comparing text masking strategies.
  • We show that word frequency distribution after masking differentiates between successful and less successful text masking strategies.
  • We carry out an in-depth comparison between existing masking strategies (CLIPA variants) and an approach that carries out masking by using frequency information directly, CLIPF. Our experiments show that CLIPF consistently outperforms syntax masking, previously considered the best choice for masking and provide a systematic view of how text masking can accelerate pre-training of VLMs while at the same time maintaining or improving VLM performance.
@@
Our work is reproducible because it uses training data sets that are openly available and small enough to allow researchers with limited computational resources to repeat our experiments, and because we release our code [^1] https://github.com/MingliangLiang3/CLIPF.git.
@@ 2.1 Vision-Language Models
Vision-language models (VLMs), such as CLIP, ALIGN, LiT, and FLIP [3, 4, 10, 8], learn image embeddings from large-scale natural language supervision using contrastive learning [14], which is important for making models transferable. Because CLIP does not release its pre-training dataset, OpenCLIP pre-trained models on several large-scale public datasets [15, 16, 17, 18] to reproduce CLIP. This implementation adheres to the scaling law observed in Neural Language Models [19] and Vision-Language Models [3], where loss decreases according to a power-law relationship with model size, dataset size, and computational resources utilized during training. While large language models (LLMs) have advanced rapidly, we focus on vision-language models (VLMs) like CLIP, as they are essential tools for bridging visual and textual modalities in current research. Although CLIP and OpenCLIP are pre-trained on very large-scale datasets, recent works like DeCLIP [20] and SLIP [5] demonstrate that Vision-language Models (VLMs) can be trained efficiently with smaller datasets, enabling researchers with limited computational resources to contribute. Moreover, masking image patches or reducing image resolution during VLM pre-training also substantially decreases computational costs while maintaining or enhancing performance, as shown by techniques like FLIP, RECLIP, CLIPA [8, 21, 12].
@@ 2.2 Text masking for VLM training
Masking text presents another viable option to speed up the pre-training of VLMs, and is particularly effective alongside strategies that employ a high ratio of image patch masking. FLIP further explored text masking techniques, namely, random and truncation masking, to enhance efficiency [8].
While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retrains the words that preceded them.
Subsequently, CLIPA [12] introduced block masking and syntax masking. As already mentioned in Section 1, CLIPA [12] shows that text masking during training can improve the performance of a VLM. CLIPA syntax masking prioritizes the retention of nouns, followed by adjectives and then other words. In [12], syntax masking outperforms the truncation, random, and block approaches. However, [12] does not vary the number of training epochs. An important contrast with our work, as previously mentioned, is our attention to the number of training epochs. Further, CLIPF, as we will see, counters the danger of giving overly high priority to nouns, leading to overfitting, and also avoids overly suppressing words of the POS category “Other”.
An initial effort to exploit frequency information for text masking during VLM training was made by SW-CLIP [22]. Unfortunately, SW-CLIPF introduces a frequency cutoff threshold, which reduces performance. In order to exploit frequency information effectively each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
3 Text Masking
In this section, we describe the text masking strategies that we study in this paper: the current state of the art, namely, the text masking strategies of CLIPA [12] (Section 3.1) and our frequency masking strategy CLIPF (Section 3.2). All these strategies accelerate pre-training by masking specific words in the text.
@@ 3 Text Masking
This function $f$ determines which words are masked to get the modified text $T_{new\_i}$. And, we define the number of input tokens as $k$. It is important to note that the number of padding tokens remains consistent across all these strategies.
3.1 Existing Text Masking Strategies (CLIPA)
The four text masking strategies from CLIPA [12], which we also investigate here, are:
@@ 3.1 Existing Text Masking Strategies (CLIPA)
3.2 Word Frequency Masking (CLIPF)
As mentioned in the introduction, subsampling of frequent words was introduced by [13] in the natural language processing (NLP) domain to accelerate the learning process for word representations. The goals are to speed up training and to address the imbalance between frequent and infrequent words in the training corpus. Our CLIPF method, uses Formula 2, which was introduced in [13], to determine the probability of masking a word during training. Thus, the complete formula to calculate the probability $P(w_{i})$ is:
$P(w_{i})=1-\sqrt{\frac{t}{f(w_{i})}}$ (2)
Here, the masking probability of the word $w_{i}$ is $P(w_{i})$, $t$ is the threshold, and $f(w_{i})$ is the frequency of the word. This formula scales the word frequency $f(w_{i})$, resulting in a much higher masking probability for frequent words than infrequent ones. It also preserves the overall ranking of word frequencies while aggressively masking high frequency words, controlled by the threshold $t$ [13].
@@ 4 Experimental Setup
Our implementation follows CLIP [3], OpenCLIP [9], FLIP [8] and CLIPA [12]. In this section, we present the details of our experimental setup.
Dataset We pre-train the models on Conceptual Captions 3M (CC3M) [24], Conceptual 12M (CC12M) [15]. These datasets were selected due to their diverse range of real-world concepts and scenes. Popular VLMs such as OpenCLIP, FLIP, DeCLIP, SLIP, A-CLIP and CLIPA [9, 8, 20, 5, 11, 12] also employ these datasets for training, leveraging their extensive and diverse content. Due to the expiration and inaccessibility of some URLs in the training set, we have downloaded 2.7 million image-text pairs for CC3M and 9.3 million image-text pairs from CC12M. The characteristics of the dataset are presented in Table 1. The CC12M dataset is 3.42 times larger than the CC3M dataset, with captions that are on average 2.15 times longer. Since the number of input tokens is fixed, the impact of different text masking ratios varies across datasets due to differences related to caption length. In an additional set of experiments, we provide results training on LAION-400M [17], a larger data set with about 300M image-text pairs.
@@
Architecture For the image encoder, we use ViT-B/16 [25] architectures with global average pooling and learnable positional embeddings. The input image size is 224 for CC3M and CC12M. For the text encoder, we use a Transformer-based model [26] and byte-pair encoding with a 49K token vocabulary. We chose these encoders because they are the largest encoders that we can use given our resources. When using full text for training, the maximum context length is 32. Zero-padding is applied to input text that is shorter than the maximum token length of the model.
We trained the models using 8 RTX A5000 GPUs with the same settings to ensure consistent conditions across all models. We experiment with different text token input sizes, namely, 32, 16, 8, 6, and 4 text tokens. As the input size gets smaller, we can increase the batch size, maximizing computational memory usage. For CC3M and CC12M, the batch sizes corresponding to the input sizes are: 664, 832, 896, 928, and 944. Note that CLIP (and therefore CLIPA and CLIPF) is trained using contrastive learning, which benefits from larger batch sizes. Note also that although we used different text masking with the same settings and batch sizes, syntax masking was slower than the other text masking strategies when conducting the experiments because POS processing for each word is time-consuming.
@@
RandomResizedCrop (50, 100) (50, 100)
Table 3: The number of image and text tokens that are processed during pre-training (CC3M and CC12M).
@@
Training Following CLIP, OpenCLIP [3, 9], we pre-train our model for 30 epochs on the CC3M and CC12M datasets. Details of the pre-training configuration are given in Table 2.
@@
During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4. To measure the training speed of CLIP, CLIPA, and CLIPF, we compare the number of text tokens processed by each model. As shown in Table 3, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios. When we pre-trained the model using image masking, the total number of tokens is 81 and the percentage of text tokens is 39.5%. When we apply 50% text masking, the total number of tokens is 65. Compared to training without text masking, this results in a speed increase of approximately 20%. Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
Fine-tuning Following FLIP, and CLIPA, in order to bridge the distribution gap between pre-training and evaluation, we fine-tuned the model without images and text masking. Note that, fine-tuning the models without masking in masking work is focused on reducing the distribution gap and applied sparingly to avoid undoing the tradeoff. The details of the fine-tuning configuration are provided in Table 2.
Table 4: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K. Reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine-tuned for an additional epoch on all data (fine-tune). “Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
@@
Evaluation setup Following CLIP [3], FLIP [8], and CLIPA [12], several classification benchmarks were used. Among these benchmarks is ImageNet-1K, a widely recognized dataset in computer vision. It is frequently used for image classification and VLM tasks and comprises 50K validation samples across 1K different classes. We filled each class into the templates provided by CLIP [3] to calculate the average of the text embeddings. We use the same evaluation settings as CLIP [3] to evaluate the other downstream datasets.
5 The Importance of Frequency Distributions
Our work is motivated by the patterns that we observe when VLMs are trained for additional epochs, shown in Fig. 1. This experiment motivates revisiting CLIPA [12], which was trained only on 6.4 epochs, in two respects. First, as already mentioned in the introduction, the conclusion that syntax masking is the best masking approach does not hold. Second, truncation masking, which looks relatively promising in low numbers of epochs, does not perform well given more epochs. In this section, we analyze the impact of masking on the training data in order to discover the reason behind this behavior.
Figure 2: Comparison of word masking probabilities for various methods. In this example, we keep four words for truncation and syntax. Numbers indicate the probability of a word being masked. Truncation keeps the first 4 words; Random and Block mask each word with a 50% probability; Syntax prioritizes retaining nouns, followed by adjectives, then others; Frequency masks words based on the frequency of the words.
We start with a basic illustration of the differences between task-masking strategies is provided in Figure 2. Truncation consistently retains the beginning of the text. With the random (random word selection) and block (random block selection) methods, each word has an equal probability of being masked, with block masking maintaining word neighborhoods. Syntax masking prioritizes the retention of nouns, followed by other part-of-speech (POS) categories.
Figure 3: The figure illustrates the distribution of top-50 words in the text before and after applying various text masking strategies. We set the text length after text masking to 6. The x-axis represents the word index, which is sorted by counts of the original data, and the y-axis shows the word frequency. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$. We remove special characters from the vocabulary.
Next, we look at how the different masking strategies impact the frequency distribution of words in the training data. Fig. 3 shows the distribution of words in the text before and after applying all of the text masking strategies. We see that in general, there is quite a large difference between the frequency of the highest-frequency words and the frequency of words with lower frequency. Such a long-tail distribution is characteristic of words occurring in text, which is known to be Zipf distributed. Most masking approaches preserve the general shape of this distribution: The frequency of a word in the masked training data tracks its frequency in the original data. The exception is syntax masking, which shows peaks and pits for specific words. Specifically, because syntax masking retains more nouns, the articles “the” and “a” are less frequent, but the nouns create spikes. The rationale for prioritizing nouns is solid. Nouns are closely related to the depicted content of images and, for this reason, are considered to be more helpful. Indeed, the helpfulness of nouns for the VLM is reflected in the fact that syntax masking learns faster than other text masking strategies. In Fig. 1, we saw that syntax masking outperforms the other masking approaches during the early epochs. However, retaining too many nouns can have two adverse effects. First, the model can overfit to certain high frequency nouns. In Fig. 3, we see that CLIPA syntax masking retains more occurrences of “person,” “illustration,” and “stock” than the other strategies. Second, retaining nouns means necessarily dropping words from other parts-of-speech (POS) categories, which may be important, as discussed in more detail below.
Considering the CLIPA truncation masking strategy, we see in Fig. 1 that the word frequency distribution of the truncated training data shows a difference with the other masking strategies. Specifically, truncation is the top line on the left of the graph, which means truncation is making the most highly frequent words in the training data relatively even more frequent. The fact that truncation impacts word frequency might seem surprising at first, but it can be explained if we consider that the probability of a word occurring changes with sentence position. With truncation, typical text beginnings, e.g., “A person is…”. remain over-represented in the data.
To dive deeper into the importance of the distribution of words over POS categories, we provide statistics for each of the different text masking strategies in Table 5. Consulting the column “Other”, we see an interesting pattern. Syntax masking changes the proportion of words in the “Other” category radically with respect to the data before masking, but truncation, random and block do not. CLIPF appears to provide a happy medium. CLIPF retains more nouns than CLIPA truncation, random, and block, but not so many that the “Other” category becomes poorly represented.
Masking NN (%) JJ (%) VB (%) OTHER (%)
Before masking 50.30% 4.98% 5.18% 39.55%
Truncation 51.72% 5.17% 5.91% 37.20%
Random 51.21% 4.82% 5.04% 38.93%
Block 49.37% 4.74% 5.24% 40.66%
Syntax 88.53% 2.69% 2.92% 5.87%
CLIPF 60.43% 5.06% 6.20% 28.24%
The importance of “Other” words is surprising because the category includes many words, which like articles, lack semantic content and are sometimes considered to have no direct connection to images. In fact, in many NLP applications, articles and other “stop words” are simply removed. However, since the CLIP text encoder is a transformer, technically it should be able to exploit the context of content words, particularly words in “Other” category. Note that with syntax masking a phrase like “smoking in the room” in the training data would always be reduced to “smoking room”, which has a different meaning. However, with the other masking approaches, “in the” would not be consistently deleted. Note specifically that block masking retains entire phrases, without always using the beginning of the sentence as truncation masking does. Also, frequency masking (CLIPF) retains entire phrases in some cases, more than would be expected with random masking.
In sum, when trained with enough epochs, overfitting effects with syntax masking occur. Our analysis has related this problem to peaks in the word frequency distribution caused by over-retention of nouns. Also, with enough epochs, truncation masking overfitting effects occur, which our analysis has related to distortion of the word frequency distribution caused by oversampling of the beginning of sentences. In sum, our analysis of word frequency distributions in the training data provides insight into why the performance of syntax and block masking falls off in Fig. 1. Further, if word frequency is indeed important, then masking that controls word-frequency directly, such as our CLIPF, can be considered to be a sensible and transparent approach.